Review Part A A 13.5 nC charge is at x 0cm and a -1.2 nC charge is at 7 cm At what point or points on the z-axis is the electric potential zero? Express your answcr using two significant figurcs. If thcrc is morc than onc answcr, give cach answer scparated by a comma. cm Submit Rcqucst AnswerReview Part A Two spherical drops of mercury each have a charge of 0.12 nC and a potential of 260 V at the surface. The two drops merge to form a single drop What is the potential at the surface of the new drop? Express your answer to two significant figurcs and include the appropriate units. V1.59 Submit Prcvious Answcrs Rcqucst Answcr X Incorrect: Try Again; 29 attempts remaining
Answer #1

Let at distance X from charge 13.5nC potential is zero

At point X

k X 13.5 x10^-9 /X = k X 1.2x10^-9/0.07- X

13.5/X= 1.2/0.07- X

1.2X + 13.5X = 0.945

14.7 X = 0.945

X = 6.4 cm

New radius R

2 × 4/3 pi X r^3 = 4/3 pi X R^3

2^1/3×r= R

V= kq/r = 260

r= 9×10^9 × 0.12× 10^-9/260= 0.42cm= 0.0042m

V = kq/r = 9× 10^9 × 2 × 0.12×10^-9/ 0.0042^1/3 = 2.16 / 0.16= 13.5V